# Label-Looping: Highly Efficient Decoding for Transducers

### v1

@@ 4.2 Results with Token-and-Duration Transducers
model name baseline ours rel. speedup
L-1 116.0 / 264.5 116.5 / 269.3 1.0X / 1.0X
L-4 243.4 / 358.1 347.7 / 632.6 1.4X / 1.8X
L-16 416.2 / 563.9 825.3 / 1615.9 2.0X / 2.9X
L-32 477.2 / 696.7 1006.8 / 2634.2 2.1X / 3.8X
XXL-1 66.1 / 241.3 66.1 / 251.7 1.0X / 1.0X
XXL-4 147.9 / 258.8 219.5 / 584.5 1.5X / 2.3X
XXL-16 243.7 / 478.3 374.8 / 1467.9 1.5X / 3.1X
XXL-32 265.2 / 616.3 394.0 / 2434.1 1.5X / 3.9X
Our TDT model experiments are shown in Table 3. Note, [11] states that TDT’s batched decoding is tricky, and proposes an approximate algorithm, which is not deterministic and requires special modifications during model training. At the time of writing, there is no open-sourced implementation of exact and general-purpose TDT batched decoding, and ours is the first such implementation available to the public. For the baseline, we adopt the approximate method from [11], which takes the minimum of predicted durations in the batch for advancement. Note: this method has non-deterministic outputs and is only included here for time comparison purposes. We can see from the Table that greater speed-ups can be seen with TDT models than conventional Transducers, and we see over 3.8X speed-up for non-encoder computation for both L and XXL models.
@@ 5 Analysis
5.1 Impact of precomputing encoder/decoder projections
In this section, we study the effects of precomputing the encoder and decoder projections, where we compare the decoding RTFx with or without the precomputation of projections. The results are in Table 4. We see that overall, precomputing the encoder and decoder outputs brings over 20% speedup for the decoding process excluding encoder computations.
@@ 5.1 Impact of precomputing encoder/decoder projections
1 147.7 179.0 1.21X
4 303.8 385.4 1.27X
32 1080.3 1403.6 1.30X
5.2 Impacts of different subsampling rates
In all experiments we report above, the encoder uses 8X subsampling. Another commonly used subsampling rate is 4X and we test our model’s speed on a publicly available model  [^3] https://catalog.ngc.nvidia.com/orgs/nvidia/teams/nemo/models/stt_en_conformer_transducer_large and show the results in Table 5. We see that with 4X models, our algorithm also significantly improves inference speed, up to 1.8X for total runtime and 2.6X for non-encoder computation.
Table 5: Performance on conformer-large models with 4x subsampling, bf16 precision. We report the total / non-encoder RTFx for decoding Librispeech-testother.
@@ 5.2 Impacts of different subsampling rates
1 48.2 / 63.2 79.5 / 130.8 1.6X / 2.1X
4 104.7 / 123.1 192.9 / 262.3 1.8X / 2.1X
16 187.2 / 243.1 341.6 / 577.5 1.8X / 2.4X
32 226.0 / 328.7 399.0 / 866.2 1.8X / 2.6X
5.3 Combining with TorchScript and CUDA Graphs
Since the speedup brought by our method is purely algorithmic, it can be combined with other methods that speed up the decoding process from other aspects such as code compilation and GPU call optimization. Here we report decoding results combining our algorithm with TorchScript [^4] https://pytorch.org/docs/stable/jit.html and with CUDA graphs [21] methods. Note, TorchScript and CUDA graph techniques aren’t compatible at the time of writing, hence we do not include experiments combining them both.
Table 6: Combining Label-looping with TorchScript or CUDA graphs. Batchsize = 32. Full-time RTFx / Non-encoder RTFx.
@@ 5.3 Combining with TorchScript and CUDA Graphs
Note, that we show this table to see that our method is compatible with other methods. Because of the scope of this paper, this is the only table with results that combine those other optimizations. The speedup reported from all other tables is solely from the algorithmic changes of the label-looping algorithm.
6 Conclusion and Future Work
In this paper, we introduce a novel set of algorithms to improve the efficiency of Transducer models. The algorithm uses a label-looping design, which minimizes the number of decoder calls, greatly reducing the runtime of decoding. Our experiments show that label-looping algorithms bring consistent speed-up over baseline batched decoding algorithms. In particular, we see a speed-up of 2.0X for batch size = 32 for conventional Transducers, and 2.1X for TDT models. Since our method is purely algorithmic, we show that it can combine with other methods like TorchScript and CUDA graphs to bring further speed up to Transducer decoding. Our algorithm will be open-sourced through the NeMo [1] toolkit.
We will continue our research efforts to improve the algorithm for future work. In particular, since our algorithm reduces the computational cost of the decoder greatly, this allows us to scale up the decoder for Transducer networks, e.g. using more layers or using more sophisticated networks like Transformers.

### v2

@@ 4.2 Results with Token-and-Duration Transducers
batch RTFx, baseline RTFx, label-looping rel. speedup
TDT-L
1 116.0 / 264.5 116.5 / 269.3 1.0 / 1.0
4 243.4 / 358.1 347.7 / 632.6 1.4 / 1.8
16 416.2 / 563.9 825.3 / 1615.9 2.0 / 2.9
32 477.2 / 696.7 1006.8 / 2634.2 2.1 / 3.8
TDT-XXL
1 66.1 / 241.3 66.1 / 251.7 1.0 / 1.0
4 147.9 / 258.8 219.5 / 584.5 1.5 / 2.3
16 243.7 / 478.3 374.8 / 1467.9 1.5 / 3.1
32 265.2 / 616.3 394.0 / 2434.1 1.5 / 3.9
Label-looping experiments for the TDT model are shown in Table 2. For the TDT baseline, we adopt the approximate method from [7], which takes the minimum of predicted durations in the batch for advancement. Note that this method has non-deterministic outputs and is only included here for time comparison purposes. Our algorithm is the first open-sourced implementation of exact and general-purpose TDT batched decoding. We observed that greater speed-ups can be seen with TDT models than conventional Transducers, and we see over 3.8X speed-up for non-encoder computation for both TDT-L and TDT-XXL models.
@@ 5 Analysis
We study the impact of precomputing the encoder and predictor projections, and compare the decoding RTFx with or without the precomputation. The results are in Table 3. We see that overall, precomputing the encoder and predictor outputs brings over 20% speedup for the decoding process, excluding encoder computations.
@@
1 147.7 179.0 1.21
4 303.8 385.4 1.27
32 1080.3 1403.6 1.30
Table 4: Performance on Conformer-Large models with 4x subsampling, bf16 precision. We report the total / non-encoder RTFx for decoding LibriSpeech test-other. WER is 3.7%.
@@
1 48.2 / 63.2 79.5 / 130.8 1.6 / 2.1
4 104.7 / 123.1 192.9 / 262.3 1.8 / 2.1
16 187.2 / 243.1 341.6 / 577.5 1.8 / 2.4
32 226.0 / 328.7 399.0 / 866.2 1.8 / 2.6
In all experiments we report above, the encoder uses 8X subsampling. Another commonly used subsampling rate is 4X, and we test our proposed algorithm on a publicly available model101010hf.co/nvidia/stt_en_conformer_transducer_large and show the results in Table 4. We see that with 4X models, our algorithm also significantly improves inference speed, up to 1.8X for total runtime and 2.6X for non-encoder computation.
It is worth mentioning that in our initial experiments, we also applied the beam search decoding algorithm to all RNNT models shown above and observed less than 0.1% absolute WER improvement while observing a several times increase in decoding time. Thus, the accelerating greedy algorithm is crucial for the optimal speed vs accuracy trade-off and opens more possibilities for scaling the models to get the best performance.
6 Combining Label-Looping with TorchScript and CUDA Graphs
Table 5: Combining Label-looping with TorchScript or CUDA graphs. Batch size = 32. Full-time RTFx / Non-encoder RTFx. Large RNNT and TDT models.
@@ 6 Combining Label-Looping with TorchScript and CUDA Graphs
Since the speedup brought by our method is purely algorithmic, it can be combined with other methods that speed up the decoding, such as code compilation and GPU call optimization. In Table 5 we report decoding results combining our algorithm with TorchScript [^11] https://pytorch.org/docs/stable/jit.html and with CUDA graphs [21] methods. TorchScript and CUDA graph techniques aren’t compatible, so, we do not include experiments combining them both.
Note that we show this table to demonstrate that our method is compatible with other methods. Because of the scope of this paper, this is the only table with results that combine those other optimizations. The speedup reported from all other tables is solely from the algorithmic changes of the label-looping algorithm.
7 Conclusion
In this paper, we introduce a novel set of algorithms to improve the efficiency of Transducer models. The algorithm uses a label-looping design, which minimizes the number of prediction network calls, greatly reducing the runtime of decoding. Our experiments show that label-looping algorithms bring consistent speed-up over baseline batched decoding algorithms. In particular, we observe a speed-up of 2.0X for batch size = 32 for conventional Transducers and 2.1X for TDT models. Since our method is purely algorithmic, we show that it can be combined with other methods, such as TorchScript and CUDA graphs, to further accelerate Transducer decoding. Our algorithm is open-sourced through the NeMo [15] toolkit.
We will continue our research efforts to improve the algorithm for future work. In particular, since our algorithm greatly reduces the computational cost of the prediction network, this allows us to scale up the predictor for Transducer networks, e.g., using more layers or more sophisticated networks like Transformers.